OpenStudy (anonymous):

What is the 6th term of the geometric sequence where a1 = 625 and a3 = 25?

Parth (parthkohli):

Gotta do it manually. \(\text{1st term} : 5^{4} \qquad 2nd:5^3 \cdots 4th:5^1\cdots 6th:{5^{-1}}\)

Parth (parthkohli):

So, what's \(5^{-1}\)?

OpenStudy (anonymous):

1/5 ?

OpenStudy (amistre64):

Jshak, do you know how to define a geometric seq?

Parth (parthkohli):

That's correct! You may answer amistre's question.

OpenStudy (anonymous):

a list of numbers that share a common ratio?

OpenStudy (amistre64):

correct :)

Parth (parthkohli):

That's very correct!

Parth (parthkohli):

Do you know the common ratio here? :)

OpenStudy (anonymous):

um 1/5?

Parth (parthkohli):

Aha!

OpenStudy (anonymous):

so the answer is 0.2 !

OpenStudy (anonymous):

thank you both!! (:

Parth (parthkohli):

You're welcome! And I didn't help you, it was just you who did the work :)

Parth (parthkohli):

Also, 0.2 is another correct notation for 1/5 ;)

OpenStudy (anonymous):

(:

OpenStudy (amistre64):

a1 = a1 a2 = a1*r a3 = a1*r*r = a1*r^2 a4 = a1*r*r*r = a1*r^3 following this pattern we can contiue a5 = a1*r^(5-1) a6 = a1*r^(6-1) an = a1*r^(n-1) we know a1 and a3 a3: 625*r^2 = 25 r^2 = 25/625 r = sqrt(25/625) r = 5/25 = 1/5 ----------------------------- to find a6 a6 = a1*r^5 = 625*(1/5)^5 = .2, yes :)